While working in the laboratory, you find a bottle containing an unknown concentration of aqueous potassium hydroxide. To determine its concentration, you take 25.0 mL of the KOH solution and titrate it with a 0.100 M solution of H2SO4. You add a bromothymol blue indicator which changes color at pH close to 7 and observe that it turns from yellow to blue when 31.2 mL of the H2SO4 solution is added. What is the concentration of the KOH solution?
Answer #1

2KOH + H2SO4 --> 2H2O + K2SO4

mmol of acid = 25*M

mmol of base = 0.1*31.2 = 3.12

2*3.12 = 25M

M = 2*3.12 /25 = 0.1704 M of KOH